A loaded grocery cart is rolling across a parking lot. You apply a constant force F=(30N)i - (40N)j to the cart as it undergoes a displacement s=(-9.0m)i - (3.0m)j. How much work does the force you apply do on the grocery cart?


1
Expert's answer
2021-05-24T09:00:44-0400

By the definition of the work done, we have:


"W=Fs,""W=[(30N)i-(40N)j]\cdot[(-9.0m)i-(3.0m)j],""W=30\ N\cdot(-9)\ m+40\ N\cdot3\ m=-150\ J."